# The Gauss map of a projective foliation

### v2

@@ 3.2. Proof of proposition 2
Note that the components of $\widehat{\mathcal{G}}$ are homogeneous polynomials of the same degree $deg(\widehat{\mathcal{G}})$. Similarly, the components of $\widehat{\Phi}$ have the same degree $deg(\widehat{\Phi})$.
The indetermination set of $\widehat{\Phi}$ has codimension $>1$, and this implies that the components of $\widehat{\Phi}$ have no common factor.
@@
In fact, from $\widehat{\Phi}\circ\widehat{\mathcal{G}}=f.I$ we get $\widehat{\Phi}\circ\widehat{\mathcal{G}}(f=0)=0$ and this implies that $\widetilde{\Phi}_{j}(\widehat{\mathcal{G}}(f=0))=0$, $1\leq j\leq n$. Since the components of $\widehat{\Phi}$ have no common components, we must have $cod(\widehat{\mathcal{G}}(f=0))\geq 2$, which implies that $cod(\mathcal{G}(V))\geq 2$.
@@
[…]
$[z]\in U\,\implies\,\widehat{\Phi}\circ\widehat{\mathcal{G}}(z)=f(z).\,z\neq 0\,\implies$
@@ 3.3. Proof of proposition 3
Let $\eta=\sum_{j=0}^{n}A_{j}(z)dz_{j}$ be an integrable form representing $\mathcal{F}$ in homogeneous coordinates, where $A_{0}$, $A_{1}$,…,$A_{n}$ are homogeneous polynomials of the same degree satisfying the relation $i_{R}\eta=\sum_{j=0}^{n}z_{j}A_{j}(z)=0$. In this case, we have
@@
Since $\widehat{\Phi}$ is birational, we get the result. ∎
@@ 3.4. Proof of theorem 2
Let $\mathcal{F}=F^{*}(\mathcal{G})$, where $F\colon\mathbb{P}^{3}\dasharrow\mathbb{P}^{2}$ has non planar fibers, except for a finite number of them. We assume that the Gauss map $G_{\mathcal{F}}$ is birational and that $G_{\mathcal{F}}$ contracts a hypersurface $S$, not $\mathcal{F}$-invariant. Remark that $S$ is not a plane, because it is not invariant. The idea is to prove that $\mathcal{F}$ has infinitely many algebraic leaves. By a theorem of Darboux, this implies that $\mathcal{F}$ has a rational first integral (see [21]). Let us prove this fact.
Since $S$ is not contracted to a point, $C:=G_{\mathcal{F}}(S)$ is a curve of ˘$\breve{\mathbb{P}}^{3}$. In particular, there is a finite subset $A\subset C$ such that for any $m\in C^{\prime}:=C\setminus A$ then $G_{\mathcal{F}}^{-1}(m)$ is a curve of $S$. Note that $G_{\mathcal{F}}^{-1}(m)$ is a plane curve because it is contracted to a point by $G_{\mathcal{F}}$. Moreover, $G_{\mathcal{F}}^{-1}(m)$ is contained in a leaf of $\mathcal{F}$, by the definition of $G_{\mathcal{F}}$.
@@
Now, since the fibers of $F$ are not plane curves $G_{\mathcal{F}}^{-1}(m)\subset F^{-1}(\hat{C})$, but $G_{\mathcal{F}}^{-1}(m)\neq F^{-1}(\hat{C})$, so that $F^{-1}(\hat{C})$ is an algebraic surface of $\mathbb{P}^{3}$. Since $\mathcal{F}=F^{*}(\mathcal{G})$ we get the leaf of $\mathcal{F}$ containing $G_{\mathcal{F}}^{-1}(m)$ is algebraic.
Now, if $m^{\prime}\in C^{\prime}$ and $m^{\prime}\neq m$ then $G_{\mathcal{F}}^{-1}(m^{\prime})\neq G_{\mathcal{F}}^{-1}(m)$, so that the leaves of $\mathcal{F}$ that contain $G_{\mathcal{F}}^{-1}(m)$ and $G_{\mathcal{F}}^{-1}(m^{\prime})$ are different. This proves the result.∎
3.5. Proofs of theorem 3 and of corollaries 2.2, 2.3 and 2.4.
@@ 3.5.1. Proof of theorem 3
2. For all $i=1,...,k$, $\Sigma\cap V_{i}$ contains exactly $deg(V_{i})$ points, say $p_{ij}$, $1\leq j\leq deg(V_{i})$.
@@
Suppose now that $\sum_{i=1}^{k}\mu(\mathcal{F},V_{i})dg(V_{i})<d^{2}+d+1$. Let $m=(d^{2}+d+1)-\sum_{i=1}^{k}\mu(\mathcal{F},V_{i}).deg(V_{i})$. In this case, if $\Sigma\simeq\mathbb{P}^{2}$ is a 2-plane of $\mathbb{P}^{3}$ transverse to $Sing_{2}(\mathcal{F})$ then the singularities of the restricted foliation $\mathcal{F}|_{\Sigma}$ are of two kinds: the points of $\Sigma\cap Sing_{2}(\mathcal{F})$ and the tangencies between the leaves of $\mathcal{F}$ and $\Sigma$. After a small perturbation of the plane $\Sigma$, we can assume that these tangencies are all of Morse type (cf. [9]). On the other hand, the number of singularities of $\mathcal{F}|_{\Sigma}$ on $\Sigma\cap\bigcup_{i}V_{i}$, counted with multiplicities, is exactly
@@
the other singularities of $\mathcal{F}|_{\Sigma}$ are the $m$ Morse tangencies. Hence $\mathcal{G}_{\mathcal{F}}^{-1}(\Sigma)$ contains $m$ points. ∎
3.5.2. Proof of corollary 2.2
First of all, let $p_{1},...,p_{N}$, $N={d^{2}+d}$, be the singularities of $\mathcal{F}_{o}|_{\Sigma}$ with generic linear part. Since $\mathcal{F}_{o}|_{\Sigma}$ has degree $d$, it must have a total of $d^{2}+d+1=N+1$ singularities, counted with multiplicity. But, the hypothesis implies that $\mu(\mathcal{F}_{o}|_{\Sigma},p_{j})=1$, $1\leq j\leq N$, so that $\mathcal{F}_{o}$ must have another singularity, say $q$, with $\mu(\mathcal{F}_{o},q)=1$. Since $\mathcal{G}_{\mathcal{F}_{o}}$ is dominant, the sigularity $q$ is necessarily a Morse center and $q$ is Morse tangency of $\mathcal{F}_{o}$ with $\Sigma$. Now, we use the known fact that singularities with non degenerate linear part are stable by small deformations of the foliation on $\mathbb{P}^{2}$ (cf [27]). This implies that:
@@
3.5.3. Proof of corollary 2.3.
Let $\mathcal{F}$ and $V_{1},...,V_{k}$ be as in the hypothesis of corollary 2.3 and $\mathcal{G}_{\mathcal{F}}\colon\mathbb{P}^{n}\dasharrow\breve{\mathbb{P}}^{n}$ be its Gauss map. The idea is to prove that a generic hyperplane of $\mathbb{P}^{n}$ has just one tangency point with the foliation. This implies that $\mathcal{G}_{\mathcal{F}}$ is birational because it is dominant.
First of all, if $\mathbb{P}^{2}\simeq\Sigma_{2}$ is a generic two plane of $\mathbb{P}^{n}$ then it has exactly one Morse tangency point with $\mathcal{F}$ at a point in $\Sigma_{2}\setminus Sing(\mathcal{F})$. This can be proved using the condition $\sum_{j=1}^{k}\mu(\mathcal{F},V_{j}).deg(V_{j})=d^{2}+d$. The proof is done in the same way as the proof of theorem 3. We leave the details for the reader.
Now, since $\mathcal{G}_{\mathcal{F}}$ is dominant a generic $(n-1)$-plane of $\mathbb{P}^{n}$ has at least one tangency of Morse type with $\mathcal{F}$ at a non singular point. Suppose by contradiction that there exists a hyperplane, say $\Sigma_{n-1}$ of $\mathbb{P}^{n}$ with at least two Morse tangencies with $\mathcal{F}$, say $m_{1}$ and $m_{2}$. In this case, let $\Sigma_{2}\subset\Sigma_{n-1}$ be a 2-plane containing the points $m_{1}$ and $m_{2}$. Then it is clear that $\Sigma_{2}$ would have Tangencies of Morse type with $\mathcal{F}$ at these points, a contradiction. ∎
3.5.4. Proof of corollary 2.4
Let $\mathcal{F}$ and $V_{1}$,…,$V_{k}$ be as before. Since $\mu(\mathcal{F},V_{i})\geq 1$ we have
@@ 3.6. Proof of theorem 4
Let $\omega$ be an integrable 1-form on $\mathbb{C}^{4}$ defining a monomial foliation on $\mathbb{P}^{3}$. We can write $\omega=\sum_{i=1}^{4}A_{i}(z)dz_{i}$, where the $A_{i^{\prime}s}$ are monomials of degree $\geq 3$.
@@
The condition $i_{R}\omega=0$ implies that $B_{1}+B_{2}+B_{3}+B_{4}=0$. Note that $B_{i}\neq 0$, $\forall i$. In fact, if $B_{4}=0$ for instance, then $\omega=A_{1}dz_{1}+A_{2}dz_{2}+A_{3}dz_{3}$ and $\mathcal{G}_{\omega}$ cannot be dominant. From now on we consider two cases for the proof:
Case 1: If $dB_{i}\wedge dB_{j}=0$ for all $i\neq j$, then $B_{i}=a_{i}B$ for some monomial $B$. By definition $z_{i}|B$ for all $i=1,2,3,4$ and since $\mathcal{G}_{\omega}$ is dominant, we conclude that $B=z_{1}z_{2}z_{3}z_{4}$, and $a_{1}+a_{2}+a_{3}+a_{4}=0$. Therefore, the foliation $\mathcal{F}_{\omega}$ has degree $2$ and it belongs to the logarithmic component $\mathcal{L}(1,1,1,1;3)$.
Case 2: There are $i\neq j$ such that $dB_{i}\wedge dB_{j}\neq 0$. We assert that after a permutation of the indexes, we can assume that $B_{1}+B_{2}=0$ and $B_{3}+B_{4}=0$. We will prove by induction on the degree $k$ of the $B_{i^{\prime}s}$ that $B_{1}=-B_{i}$ for some $i>1$.
For $k=1$ the proof is left to the reader. Suppose by induction that the assertion is true for $k\geq 1$ and let us prove that it is true for $k+1$.
Since $B_{1}\neq 0$ there exists $j$ such that $\frac{\partial B_{1}}{\partial z_{j}}\neq 0$ and $\sum_{i=1}^{4}\frac{\partial B_{i}}{\partial z_{j}}=0$. If $\frac{\partial B_{i}}{\partial z_{j}}\neq 0$ for all $i>1$, then $\frac{\partial B_{i}}{\partial z_{j}}$ is a monomial of degree $k$ for all $i$ and the induction hypothesis implies that there is $l>1$ such that $\frac{\partial B_{1}}{\partial z_{j}}=-\frac{\partial B_{l}}{\partial z_{j}}$, so that $B_{1}+B_{l}=0$, because they are monomials. After a permutation of the indexes, we can assume that $B_{1}+B_{2}=0$ and $B_{3}+B_{4}=0$.
If there is $i>l>1$ such that $\frac{\partial B_{i}}{\partial z_{j}}=\frac{\partial B_{l}}{\partial z_{j}}=0$, then after a permutation of the indexes, we can assume that $\frac{\partial B_{1}}{\partial z_{j}}=-\frac{\partial B_{2}}{\partial z_{j}}\neq 0$, which implies $B_{1}+B_{2}=0$ and $B_{3}+B_{4}=0$.
The remaining possility is that for just one $i>1$, for instance $i=4$, we have $\frac{\partial B_{4}}{\partial z_{j}}=0$. But, as the reader can check, this is not possible.
@@
Now, if $B_{1}=-B_{2}$ and $B_{3}=-B_{4}$ we have $\omega=B_{1}\Theta_{12}+B_{3}\Theta_{34}$, where $\Theta_{ij}=\frac{dz_{i}}{z_{i}}-\frac{dz_{j}}{z_{j}}$. Since $d\Theta_{ij}=0$ and $\Theta_{ij}=\frac{z_{j}}{z_{i}}d\left(\frac{z_{i}}{z_{j}}\right)$, the reader can check that integrability condition $\omega\wedge d\omega=0$ can be written as
$00=d\left(\frac{B_{1}}{B_{3}}\right)\wedge\Theta_{12}\wedge\Theta_{34}=0\,\implies\,d\left(\frac{B_{1}}{B_{3}}\right)\wedge d\left(\frac{z_{1}}{z_{2}}\right)\wedge d\left(\frac{z_{3}}{z_{4}}\right)=$ 0
$\implies\,\,\frac{B_{1}}{B_{3}}=\lambda\left(\frac{z_{1}}{z_{2}}\right)^{\pm p}\left(\frac{z_{3}}{z_{4}}\right)^{\pm q}\,,$
@@
where $p,q\in\mathbb{Z}_{\geq 0}$ and $p+q\geq 1$. In particular, $B_{1}/B_{3}$ is non-constant and after a linear change of variables, we can assume $\lambda=1$. This implies,
$B_{3}^{-1}.\,\omega=\pm\frac{B_{1}}{B_{3}}\left(\frac{dz_{1}}{z_{1}}-\frac{dz_{2}}{z_{2}}\right)+\left(\frac{dz_{3}}{z_{3}}-\frac{dz_{4}}{z_{4}}\right)=$
$=B_{3}^{-1}.\,\omega=\left(\frac{z_{1}}{z_{2}}\right)^{\pm p}\left(\frac{z_{3}}{z_{4}}\right)^{\pm q}\left[\pm\frac{z_{2}}{z_{1}}d\left(\frac{z_{1}}{z_{2}}\right)+\frac{z_{4}}{z_{3}}d\left(\frac{z_{3}}{z_{4}}\right)\right]$
@@
3.7. Proofs of proposition 4 and of corollary 2.5

### v3

@@ 3.2. Proof of Proposition 2
[…]
The indetermination set of $\widehat{\Phi}$ has codimension >1>1, and this implies that the components of $\widehat{\Phi}$ have no common factor.
@@
In fact, from $\widehat{\Phi}\circ\widehat{\mathcal{G}}=f.I$ we get $\widehat{\Phi}\circ\widehat{\mathcal{G}}(f=0)=0$ and this implies that $\widehat{\Phi}_{j}(\widehat{\mathcal{G}}(f=0))=0$, $1\leq j\leq n$. Since the components of $\widehat{\Phi}$ have no common components, we must have $cod(\widehat{\mathcal{G}}(f=0))\geq 2$, which implies that $cod(\mathcal{G}(V))\geq 2$.
@@
has dimension n1n-1 and is Zariski dense in $(h=0)$. Let $z]\in U$. In homogeneous coordinates we have
$z]\in U\,\implies\,\widehat{\Phi}\circ\widehat{\mathcal{G}}(z)=f(z).\,z\neq 0\,\implies$
@@ 3.3. Proof of Proposition 3
Let $\eta=\sum_{j=0}^{n}A_{j}(z)dz_{j}$ be an integrable form representing $\mathcal{F}$ in homogeneous coordinates, where $A_{0}$, $A_{1}$$A_{n}$ are homogeneous polynomials of the same degree satisfying the relation $i_{R}\eta=\sum_{j=0}^{n}z_{j}A_{j}(z)=0$. In this case, we have
@@
Since $\widehat{\Phi}$ is birational, we obtain $\mathcal{G}_{(\mathcal{G})_{*}(\mathcal{F})}=\mathcal{G}_{\mathcal{F}}^{-1}$. ∎
@@ 3.4. Proof of Theorem 2
Let $\mathcal{F}=F^{*}(\mathcal{G})$, where $F\colon\mathbb{P}^{3}\dasharrow\mathbb{P}^{2}$ has non planar fibers, except for a finite number of them. We assume that the Gauss map $G_{\mathcal{F}}$ is birational and that $G_{\mathcal{F}}$ contracts a hypersurface $S$, not $\mathcal{F}$-invariant. Remark that $S$ is not a plane because, using the condition that $G_{\mathcal{F}}(S)$ is a point, we would have that $S$ is invariant. Then, the idea is to prove that $\mathcal{F}$ has infinitely many algebraic leaves. By a theorem of Darboux, this implies that $\mathcal{F}$ has a rational first integral (see [23]). Let us prove this fact.
Since $S$ is not contracted to a point, $C:=G_{\mathcal{F}}(S)$ is a curve of $\breve{\mathbb{P}}^{3}$. In particular, there is a finite subset $A\subset C$ such that for any $m\in C^{\prime}:=C\setminus A$, $G_{\mathcal{F}}^{-1}(m)$ is a curve of $S$. Note that $G_{\mathcal{F}}^{-1}(m)$ is a plane curve because it is contracted to a point by $G_{\mathcal{F}}$. Moreover, $G_{\mathcal{F}}^{-1}(m)$ is contained in a leaf of $\mathcal{F}$, by the definition of $G_{\mathcal{F}}$.
@@
$(\mathbb{P}^{2},\widehat{C})\underset{F}{\longleftarrow}(\mathbb{P}^{3},S\supset G_{\mathcal{F}}^{-1}(m))\underset{G_{\mathcal{F}}}{\longrightarrow}(\breve{\mathbb{P}}^{3},C\ni m)$
@@
Now, if $m^{\prime}\in C^{\prime}$ and $m^{\prime}\neq m$ then $G_{\mathcal{F}}^{-1}(m^{\prime})\neq G_{\mathcal{F}}^{-1}(m)$, because these curves are contracted to a point by $G_{\mathcal{F}}$. The intersections $G_{\mathcal{F}}^{-1}(m)\cap S$ define a foliation of the surface $S$ by algebraic curves. If $\mathcal{L}$ is a leaf of $\mathcal{F}$ containing $G_{\mathcal{F}}^{-1}(m)$, $\mathcal{L}$ intersects $S$ in at most a countable number (in fact finite, since it is algebraic) of connected components. Therefore, for generic $m$ and $m^{\prime}$, the corresponding leaves of $\mathcal{F}$ are distinct. Consequently, $\mathcal{F}$ has infinitely many algebraic leaves. This proves the result.∎
3.5. Proofs of Theorem 3 and of Corollaries 2, 3 and 5.
@@ 3.5.1. Proof of Theorem 3
2. For all i=1,,ki=1,...,k, $\Sigma\cap V_{i}$ contains exactly $deg(V_{i})$ points, say $p_{ij}$, $1\leq j\leq deg(V_{i})$.
@@
Let $m=(d^{2}+d+1)-\sum_{i=1}^{k}\mu(\mathcal{F},V_{i}).deg(V_{i})$. In this case, if $\Sigma\simeq\mathbb{P}^{2}$ is a 2-plane of $\mathbb{P}^{3}$ transverse to $Sing_{2}(\mathcal{F})$ then the singularities of the restricted foliation $\mathcal{F}|_{\Sigma}$ are of two kinds: the points of $\Sigma\cap Sing_{2}(\mathcal{F})$ and the tangencies between the leaves of $\mathcal{F}$ and $\Sigma$. After a small perturbation of the plane $\Sigma$, we can assume that these tangencies are all of Morse type (see [10]). On the other hand, the number of singularities of $\mathcal{F}|_{\Sigma}$ in $\Sigma\cap\bigcup_{i}V_{i}$, counted with multiplicities, is exactly
@@
the other singularities of $\mathcal{F}|_{\Sigma}$ are the $m$ Morse tangencies (note that $m$ could be zero). Hence $\mathcal{G}_{\mathcal{F}}^{-1}(\Sigma)$ contains $m$ points. ∎
3.5.2. Proof of corollary 2
First of all, let $p_{1},...,p_{N}$, $N={d^{2}+d}$, be the singularities of $\mathcal{F}_{o}|_{\Sigma}$ with generic linear part. Since $\mathcal{F}_{o}|_{\Sigma}$ has degree $d$, it must have a total of $d^{2}+d+1=N+1$ singularities, counted with multiplicity. However, the hypothesis implies that $\mu(\mathcal{F}_{o}|_{\Sigma},p_{j})=1$, $1\leq j\leq N$, so that $\mathcal{F}_{o}$ must have another singularity, say $q$, with $\mu(\mathcal{F}_{o},q)=1$. Since $\mathcal{G}_{\mathcal{F}_{o}}$ is dominant, the singularity $q$ is necessarily a Morse center and $q$ is a Morse tangency of $\mathcal{F}_{o}$ with $\Sigma$. Now, we use the known fact that singularities with non degenerate linear part are stable by small deformations of the foliation on $\mathbb{P}^{2}$ (see section 2.4 and corollary 2.7 of [28]). This implies that:
@@ 3.5.2. Proof of corollary 2
3.5.3. Proof of corollary 3.
Let $\mathcal{F}$ and $V_{1},...,V_{k}$ be as in the hypothesis of Corollary 3 and let $\mathcal{G}_{\mathcal{F}}\colon\mathbb{P}^{n}\dasharrow\breve{\mathbb{P}}^{n}$ be its Gauss map. The idea is to prove that a generic hyperplane of $\mathbb{P}^{n}$ has just one tangency point with the foliation. This implies that $\mathcal{G}_{\mathcal{F}}$ is birational.
First of all, if $\mathbb{P}^{2}\simeq\Sigma_{2}$ is a generic two-plane of $\mathbb{P}^{n}$ then it has exactly one Morse tangency point with $\mathcal{F}$ at a point in $\Sigma_{2}\setminus Sing(\mathcal{F})$. This can be proved using the condition $\sum_{j=1}^{k}\mu(\mathcal{F},V_{j}).deg(V_{j})=d^{2}+d$. The proof is done in the same way as the proof of Theorem 3. We leave the details to the reader.
Now, since $\mathcal{G}_{\mathcal{F}}$ is dominant, a generic (n1)(n-1)-plane of $\mathbb{P}^{n}$ has at least one tangency of Morse type with $\mathcal{F}$ at a non singular point. Suppose by contradiction that there exists a hyperplane, say $\Sigma_{n-1}$ of $\mathbb{P}^{n}$ with at least two Morse tangencies with $\mathcal{F}$, say $m_{1}$ and $m_{2}$. In this case, let $\Sigma_{2}\subset\Sigma_{n-1}$ be a 2-plane containing the points $m_{1}$ and $m_{2}$. Then it is clear that for a generic choice of such a 2-plane, $\Sigma_{2}$ would have tangencies of Morse type with $\mathcal{F}$ at these points, a contradiction. ∎
3.5.4. Proof of corollary 4
Since $\mathbb{P}^{m}$ is linearly embedded in $\mathbb{P}^{n}$, the components of the singular set of the restriction $\mathcal{F}_{|\mathbb{P}^{m}}$ satisfy the formula of the previous corollary, so we have the result. ∎
3.5.5. Proof of corollary 5
Let $\mathcal{F}$ and $V_{1}$$V_{k}$ be as before. Since $\mu(\mathcal{F},V_{i})\geq 1$ we have
@@ 3.6. Proof of Theorem 4
Let $\omega$ be an integrable 1-form on $\mathbb{C}^{4}$ defining a monomial foliation on $\mathbb{P}^{3}$ of degree $d\geq 2$. We can write $\omega=\sum_{i=1}^{4}A_{i}(z)dz_{i}$, where $A_{i^{\prime}s}$ are monomials of degree $\geq 3$. Let
@@
The condition $i_{R}\omega=0$ implies that $B_{1}+B_{2}+B_{3}+B_{4}=0$. Note that $B_{i}\neq 0$, for all $i$. In fact, if $B_{4}=0$ for instance, then $\omega=A_{1}dz_{1}+A_{2}dz_{2}+A_{3}dz_{3}$ and $\mathcal{G}_{\omega}$ cannot be dominant. We also ask that the monomials $A_{1}$, $A_{2}$, $A_{3}$, $A_{4}$ have no variables in common to have codimension of the singular set $\geq 2$.
Observe that each $B_{i}$ is a monomial of degree $d+2$, with exponent vector $(a_{11}+1,a_{12},a_{13},a_{14})$, $(a_{11},a_{12}+1,a_{13},a_{14})$, $(a_{11},a_{12},a_{13}+1,a_{14})$, $(a_{11},a_{12},a_{13},a_{14}+1)\in\mathbb{Z}^{4}_{\geq 0}$, for $i=1,2,3,4$; respectively. Since the sum $B_{1}+B_{2}+B_{3}+B_{4}$ vanishes identically, the monomials must be grouped according to identical exponent vectors, with the coefficients in each group summing to zero. With this observation, only the following cases can occur.
Case 1: All monomials in $\{B_{1},B_{2},B_{3},B_{4}\}$ have exactly the same vector of exponents in $\mathbb{Z}^{4}_{\geq 0}$. In this case and since $A_{i}\textasciiacute s$ do not have variables in common, we have $B_{i}=\lambda_{i}z_{1}z_{2}z_{3}z_{4}$ for all $i=1,2,3,4$ and $\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}=0$. Therefore, the foliation $\mathcal{F}_{\omega}$ has degree $2$ and belongs to the logarithmic component $\mathcal{L}(1,1,1,1;3)$.
Case 2: If the exponent vectors are not all equal, there cannot exist a single exponent vector differing from the others, because in this case the sum of the monomials is not zero. Therefore, with 4 exponent vectors, we only have the following possibility: the set $\{B_{1},B_{2},B_{3},B_{4}\}$ splits into exactly two disjoint subsets, each containing two monomials with identical exponents. And the sum of their coefficients is zero.
@@
Then, after a permutation of the indexes, we can assume that $B_{1}+B_{2}=0$ and $B_{3}+B_{4}=0$. Then we have $\omega=B_{1}\Theta_{12}+B_{3}\Theta_{34}$, where $\Theta_{ij}=\frac{dz_{i}}{z_{i}}-\frac{dz_{j}}{z_{j}}$. Since $d\Theta_{ij}=0$ and $\Theta_{ij}=\frac{z_{j}}{z_{i}}d\left(\frac{z_{i}}{z_{j}}\right)$, the reader can check that integrability condition $\omega\wedge d\omega=0$ can be written as
$0=d\left(\frac{B_{1}}{B_{3}}\right)\wedge\Theta_{12}\wedge\Theta_{34}=0\,\implies\,d\left(\frac{B_{1}}{B_{3}}\right)\wedge d\left(\frac{z_{1}}{z_{2}}\right)\wedge d\left(\frac{z_{3}}{z_{4}}\right)=0.$
With this we conclude that equations $a_{11}+1-a_{31}=a_{32}-a_{12}=p$, $a_{13}-a_{33}-1=a_{34}-a_{14}=q$ are satisfied, to have
$\frac{B_{1}}{B_{3}}=\frac{\lambda_{1}}{\lambda_{3}}\left(\frac{z_{1}}{z_{2}}\right)^{p}\left(\frac{z_{3}}{z_{4}}\right)^{q}\,$,
where $\lambda\in\mathbb{C}^{*}$. Since $z_{2}|B_{1}$ and $z_{4}|B_{3}$, then $a_{12}>0$ and $a_{34}>0$, this implies that $a_{32}=0$ and $a_{14}=0$; otherwise $z_{2}$ or $z_{4}$ divides $\omega$. Similarly, we can see that if $a_{11}>0$ or $a_{33}>0$ then $z_{1}|A_{1},A_{2}$ or $z_{3}|A_{3},A_{4}$, this implies $a_{31}=0$ or $a_{13}=0$, in the first case $p=-a_{12}=a_{11}+1>0$ and, in the second case, $q=a_{34}=-a_{33}-1<0$. Both produce a contradiction; therefore, $a_{11}=a_{33}=0$. With this we obtain the following specific 1-form that defines the foliation:
$\omega=\lambda_{1}z_{1}z_{2}^{a_{12}}z_{3}^{a_{13}}\Theta_{12}+\lambda_{3}z_{1}^{1+a_{12}}z_{3}z_{4}^{a_{13}-1}\Theta_{34},$
@@
where $a_{12}\geq 1$ and $a_{13}\geq 2$. Moreover, we have $p=-a_{12}=1-a_{31}<0$ and $q=a_{34}=a_{13}-1>0$. After a linear change of variables, we can assume $\frac{\lambda_{1}}{\lambda_{3}}=1$. With all this, we finally conclude that
$B_{3}^{-1}.\,\omega=\frac{B_{1}}{B_{3}}\left(\frac{dz_{1}}{z_{1}}-\frac{dz_{2}}{z_{2}}\right)+\left(\frac{dz_{3}}{z_{3}}-\frac{dz_{4}}{z_{4}}\right)=$
$\left(\frac{z_{1}}{z_{2}}\right)^{p}\left(\frac{z_{3}}{z_{4}}\right)^{q}\left[\frac{z_{2}}{z_{1}}d\left(\frac{z_{1}}{z_{2}}\right)\right]+\frac{z_{4}}{z_{3}}d\left(\frac{z_{3}}{z_{4}}\right)$
@@
3.7. Proofs of Proposition 4 and of Corollary 6